# Language Modeling on a SpiNNaker 2 Neuromorphic Chip

### 14 Dec 2023 v1

@@ Abstract
As large language models continue to scale in size rapidly, so too does the computational power required to run them. Event-based networks on neuromorphic devices offer a potential way to reduce energy consumption for inference significantly. However, to date, most event-based networks that can run on neuromorphic hardware, including spiking neural networks (SNNs), have not achieved task performance even on par with LSTM models for language modeling. As a result, language modeling on neuromorphic devices has seemed a distant prospect. In this work, we demonstrate the first-ever implementation of a language model on a neuromorphic device – specifically the SpiNNaker 2 chip – based on a recently published event-based architecture called the EGRU. SpiNNaker 2 is a many-core neuromorphic chip designed for large-scale asynchronous processing, while the EGRU is architected to leverage such hardware efficiently while maintaining competitive task performance. This implementation marks the first time a neuromorphic language model matches LSTMs, setting the stage for taking task performance to the level of large language models. We also demonstrate results on a gesture recognition task based on inputs from a DVS camera. Overall, our results showcase the feasibility of this neuro-inspired neural network in hardware, highlighting significant gains versus conventional hardware in energy efficiency for the common use case of single batch inference.
@@ VI Outlook
The successful implementation of a language model on the SpiNNaker 2 chip using the EGRU event-based architecture represents a significant milestone in the field of neuromorphic computing. We compare our implementation with the one on the Nvidia GPU and show a real energy advantage in the single batch size setting, which is expected to be the most relevant for inference, especially on edge devices. We also identified several bottlenecks in our implementation that need improvement for even further efficiency. In particular, quantizing the model will allow us to work with even tighter memory constraints to fit larger models and further increase energy efficiency, albeit with a slight reduction in task performance. Quantization will also allow us to take advantage of the MAC accelerator available on the Spinnaker 2 chip. We also plan to further scale the deployed model to multiple chips, as Spinnaker 2 is designed for efficient distributed computing.
@@
Acknowledgment
Khaleel Khan is Funded by the German Federal Ministry of Education and Research (BMBF), funding reference 16ME0729K, joint project ”EVENTS”. Mark Schöne is fully funded by the Bosch Research Foundation. David Kappel is funded by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) project ESCADE (01MN23004A).

### 17 Jan 2024 v2

@@ Abstract
As large language models continue to scale in size rapidly, so too does the computational power required to run them. Event-based networks on neuromorphic devices offer a potential way to reduce energy consumption for inference significantly. However, to date, most event-based networks that can run on neuromorphic hardware, including spiking neural networks (SNNs), have not achieved task performance even on par with LSTM models for language modeling. As a result, language modeling on neuromorphic devices has seemed a distant prospect. In this work, we demonstrate the first-ever implementation of a language model on a neuromorphic device – specifically the SpiNNaker2 chip – based on a recently published event-based architecture called the EGRU. SpiNNaker2 is a many-core neuromorphic chip designed for large-scale asynchronous processing, and the EGRU is architected to leverage such hardware efficiently while maintaining competitive task performance. This implementation marks the first time a neuromorphic language model matches LSTMs, setting the stage for taking task performance to the level of large language models. We also demonstrate results on a gesture recognition task based on inputs from a DVS camera. Overall, our results showcase the feasibility of this neuro-inspired neural network in hardware, highlighting significant gains versus conventional hardware in energy efficiency for the common use case of single batch inference.
@@ VI Outlook
The successful implementation of a language model on the SpiNNaker2 chip using the EGRU event-based architecture represents a significant milestone in the field of neuromorphic computing. We compare our implementation with the one on the Nvidia GPU and show a real energy advantage in the single batch size setting, which we expect to be the most relevant for inference, especially on edge devices. We also identified several bottlenecks in our implementation that need improvement for even further efficiency. In particular, quantizing the model will allow us to work with even tighter memory constraints to fit larger models and further increase energy efficiency, albeit with a slight reduction in task performance. Quantization will also allow us to take advantage of the MAC accelerator available on the SpiNNaker2 chip. We also plan to further scale the deployed model to multiple chips, as SpiNNaker2 is designed for efficient distributed computing.